# Equi-RO: A 4D mmWave Radar Odometry via Equivariant Networks

### v1

@@ III-C Graph Construction
After point cloud pre-processing, the graph is constructed by treating each point as a node. A cost function $f(i,j)$ is defined to select node pairs for edge generation, preferring neighboring points with similar velocities:
@@
where $\lambda$ is a weighting parameter. An edge is established between nodes $i$ and $j$ if $f(i,j)$ is below the threshold $F_{t}$.
Subsequently, node and edge features are specified, including the node equivariant feature $\mathbf{X}\in\mathbb{R}^{N\times 3}$, the node invariant feature $\mathbf{Y}\in\mathbb{R}^{N\times 3}$, and the edge invariant feature $\mathbf{Z}\in\mathbb{R}^{E\times 2}$, where $E$ is the total number of edges:
$\mathbf{X}$ $=\{\mathbf{X}_{i}\}=\{\bm{\rho}_{i}\},$ (7)
$\mathbf{Y}$ $=\{\mathbf{Y}_{i}\}=\{[v_{i}^{dop^{{}^{\prime}}},r_{i},d_{i}]\},$ (8)
$\mathbf{Z}$ $=\{\mathbf{Z}_{ij}\}=\{[v_{ij},\left\|\bm{\rho}_{i}-\bm{\rho}_{j}\right\|]\},j\in\mathcal{N}(i),$ (9)
where i=1,2,,Ni=1,2,...,N, $r_{i}$ denotes the RCS value of point $i$, $\mathcal{N}(i)$ is the set of neighbors of $i$ in the graph, and $d_{i}$ is the degree of node $i$, i.e., the number of edges connected to it.
The SE(3)-equivariance of $\mathbf{X}$ is straightforward, as it represents the Euclidean coordinates of the points. The SO(3)-invariance of $v_{i}^{dop^{{}^{\prime}}}$ and the SE(3)-invariance of $r_{i}$ and $d_{i}$ make the node feature $\mathbf{Y}$ invariant. Likewise, in the edge feature $\mathbf{Z}$, the compensated edge relative velocity $v_{ij}$ remains SE(3)-invariant for nodes with the same velocity, which can be reinforced by properly defining the cost function $f(i,j)$. The Euclidean distance $\left\|\bm{\rho}_{i}-\bm{\rho}_{j}\right\|$ between nodes $i$ and $j$ is also SE(3)-invariant.
@@ III-D Network Design
Following graph construction and feature preparation, we establish a novel framework that separately processes the input equivariant and invariant features. The framework employs SE(3)-equivariant networks to handle different feature types, and estimates the relative transformation between the two frames based on the extracted final features.
@@ III-D1 Feature Extraction
We employ two main architectures for feature extraction. The first one is an SE(3)-equivariant Dynamic Graph Convolutional Neural Network (DGCNN) [23] processing the SE(3)-equivariant node feature $\mathbf{X}$ of two consecutive point clouds and preserving its equivariance. The second one is an SE(3)-equivariant GNN (EGNN) module that, by integrating the processed node feature $\mathbf{X^{\prime}}$ with the node and edge invariant feature $\mathbf{Y}$ and $\mathbf{Z}$, outputs the final feature $\mathbf{F}$.
The first DGCNN network dynamically aggregates features from each point and its neighbors by a local graph, thereby generating richer local feature representations. Since SE(3)-equivariance is composed of translational and rotational equivariance in 3D space, we first achieve translational equivariance by translating the earlier point cloud frame to the origin of the current frame using the estimated ego-velocity and timestamp difference. This allows us to focus on rotational equivariance (i.e. SO(3)-equivariance), which is handled by a stack of LN [20] network layers in the DGCNN. LN modifies traditional MLP layers such as Linear, ReLU and Pooling layers for equivariance on semisimple Lie groups, and helps to preserve SO(3)-equivariance in our case. After three iterations, the network outputs the processed node feature $\mathbf{X^{\prime}}$, whose dimension is lifted to $N\times(C^{\prime}\times 3)$ by LN and retains SE(3)-equivariance.
The processed node feature $\mathbf{X^{\prime}}$ is then passed to the second EGNN module together with $\mathbf{Y}$ and $\mathbf{Z}$ for further graph feature aggregation. For each node $i$ and its neighbors $\mathcal{N}(i)$ connected by edges, the equivariant and invariant features are aggregated and updated as follows:
@@
$\mathbf{X}^{{}^{\prime}}_{i}$ $=\mathbf{X}^{{}^{\prime}}_{i}+C\underset{j\in\mathcal{N}(i)}{\sum}(\mathbf{X}^{{}^{\prime}}_{i}-\mathbf{X}^{{}^{\prime}}_{j})\phi_{E}(\mathbf{m}_{ij}),$ (12)
where $\mathbf{m}_{ij}$ represents the invariant message passed from node $j$ to node $i$ under the Message Passing Neural Network (MPNN) framework [24], $\mathbf{X}^{{}^{\prime}}_{i}$ and $\mathbf{Y}_{i}$ are processed node equivariant and invariant feature of node $i$, respectively. $C$ is a constant defined as the number of elements in set $\mathcal{N}(i)$, and $\phi_{I_{1}},\phi_{I_{2}},\phi_{E}$ are MLP networks.
Differing from the original EGNN that uses relative squared distances $\left\|\mathbf{X}^{{}^{\prime}}_{i}-\mathbf{X}^{{}^{\prime}}_{j}\right\|^{2}$ for invariant feature extraction, inspired by [25], we employ a Gaussian kernel function $\exp(\gamma\left\|\mathbf{X}_{i}^{{}^{\prime}}-\mathbf{X}_{j}^{{}^{\prime}}\right\|)$ with a learnable parameter $\gamma$ to transform the SE(3)-equivariant features $\mathbf{X}$ into SE(3)-invariant representations.
After four iterations, the network outputs the final node equivariant feature set $\mathbf{X}_{F}=\{\mathbf{X}^{{}^{\prime}}_{i}\}$ and final node invariant feature set $\mathbf{Y}_{F}=\{\mathbf{Y}_{i}\}$ for i=1,2,,Ni=1,2,...,N. The final feature $\mathbf{F}$ is defined as the concatenation of $\mathbf{X}_{F}$ and $\mathbf{Y}_{F}$.
@@ III-D2 Transformation Estimation
In this module, the $\mathbf{L}_{2}$ distances between the final features $\mathbf{F}$ of points in the two point cloud frames are computed. The $M$ pairs with the smallest feature distances are selected as keypoint correspondences. A learnable similarity metric is then defined as:
$s_{ij}=\exp(-\beta(d_{ij}-\alpha)),$ (13)
where $s_{ij}$ denotes the similarity between point $i$ in the first frame and point $j$ in the second frame, $d_{ij}$ is their feature distance, and $\alpha,\beta$ are learnable parameters that enhance the robustness of the pipeline to outliers.
Given the similarity matrix $\mathbf{S}=\{s_{ij}\}_{M\times M}$, we apply the differentiable Sinkhorn normalization algorithm [26] to obtain a soft assignment of correspondences. After $K$ iterations, the assigning weight matrix $\mathbf{W}=\{w_{ij}\}_{M\times M}$ is obtained. A weighted Singular Value Decomposition (SVD) is then used to estimate the relative transformation $(\mathbf{R},\mathbf{t})$ between the two frames.
@@ III-D3 Loss Function
To supervise the transformation estimation, we design a novel loss function that jointly considers rotation, translation, and orientation errors. The estimated rotation $\mathbf{R}$ is first converted to Euler angles $\mathbf{E}$, and the loss $\mathcal{L}$ between $(\mathbf{E},\mathbf{t})$ and the ground truth $(\mathbf{E}_{gt},\mathbf{t}_{gt})$ consists of four terms: rotation loss $\mathcal{L}_{r}$, translation loss $\mathcal{L}_{t}$, pitch loss $\mathcal{L}_{p}$, and yaw loss $\mathcal{L}_{y}$:
$\mathcal{L}_{r}$ $=\left\|\mathbf{E}-\mathbf{E}_{gt}\right\|,$ (14)
$\mathcal{L}_{t}$ $=\left\|\mathbf{t}-\mathbf{t}_{gt}\right\|,$ (15)
$\mathcal{L}_{p}$ $=\left\|\mathbf{E}^{pitch}-\mathbf{E}_{gt}^{pitch}\right\|,$ (16)
$\mathcal{L}_{y}$ $=\left\|\mathbf{E}^{yaw}-\mathbf{E}_{gt}^{yaw}\right\|,$ (17)
where $\mathbf{E}^{pitch}$ and $\mathbf{E}_{gt}^{pitch}$ denote the pitch angles, and $\mathbf{E}^{yaw}$ and $\mathbf{E}_{gt}^{yaw}$ denote the yaw angles, in the estimated and ground truth rotations, respectively. The pitch and yaw losses are included to compensate for the relatively low elevation and azimuth resolution of 4D mmWave radar measurements.
@@ IV-A Datasets
Two datasets are used for evaluation: the publicly available NTU4DRadLM dataset [6] and a self-collected dataset acquired using a vehicle-mounted 4D mmWave radar and high-precision GPS.
@@ IV-A1 NTU4DRadLM Dataset
The NTU4DRadLM dataset is specifically designed for 4D radar-based localization and mapping. It provides accurate ground-truth odometry and includes sequences from both low-speed robot platforms and high-speed unmanned vehicles. The dataset was collected using the Oculii Eagle 4D radar, featuring an azimuth resolution of $0.5^{\circ}$, an elevation resolution of $1^{\circ}$, and an operating frequency of 12 Hz. NTU4DRadLM contains six splits. We use loop1 for training, loop2 for validation, and the remaining four splits for testing.
@@ IV-A2 Self-Collected Dataset
To further assess generalization, we constructed a campus dataset using the Geometrical Pal R7861B 4D radar. As shown in Fig. 4, the radar was mounted on an SUV equipped with a high-precision GPS system to provide ground-truth odometry. The radar has an azimuth resolution of $1.9^{\circ}$, an elevation resolution of $3.5^{\circ}$, and operates at 15 Hz. For evaluation, models of all learning-based algorithms trained on the loop1 split of NTU4DRadLM are directly tested on this dataset to compare cross-dataset generalization.
@@
IV-B Experimental Setup
IV-B1 Training Details
The proposed model is trained on a single NVIDIA A100 GPU using the Adam optimizer with a weight decay of $10^{-3}$ for 50 epochs , taking approximately 6 hours. The initial learning rate is set to $10^{-4}$ and decays by a factor of 0.1 every 10 epochs. The batch size is set at 4.
IV-B2 Compared algorithms
For baseline comparison, we select several traditional point cloud registration methods, including Iterative Closest Point (ICP) [27], Normal Distributions Transform (NDT) [28], and Generalized ICP (GICP) [29]. In addition, we evaluate the State-of-the-Art (SOTA) 4D mmWave radar odometry method Adaptive Probability Distribution-GICP (APDGICP), which serves as the front end of the 4DRadarSLAM framework [2].
For learning-based methods, we include the self-supervised odometry method RaFlow [13] and the recent supervised method CAO-RONet [15].
IV-B3 Metrics
We adopt the evo library [30] to compute the relative translation error ($t_{rel}$) and relative rotation error ($r_{rel}$) for each algorithm, by comparing registered trajectories with the ground-truth.
Figure 5: Projection of adjacent point clouds onto the corresponding images, where point colors represent compensated node velocities and white circles denote selected keypoints used for matching.
Figure 6: Comparative odometry results of Equi-RO and baseline algorithms on the NTU4DRadLM dataset for the cp (a), garden (b), nyl (c), and loop3 (d) splits. Regions with large rotational motion are highlighted with magnified insets for clarity.
Figure 7: Translation error (a) and rotation error (b) profiles of Equi-RO and selected baseline algorithms over the full cp split trajectory.
Figure 8: Comparative odometry results of Equi-RO and baseline algorithms on the self-collected dataset.
@@ IV-B3 Metrics
IV-C Experimental Results
Table I presents a quantitative comparison of our Equi-RO against several baseline algorithms. In NTU4DRadLM dataset, our method achieves the lowest translation and rotation errors on the cp, garden, and loop3 splits, and ranks second best on the nyl split. Traditional methods such as GICP and APDGICP attain competitive results, while learning-based approaches RaFlow and CAO-RONet exhibit inferior performance. This discrepancy is likely due to their design and evaluation primarily on the earlier VoD dataset [31], leading to limited generalization on the newer NTU4DRadLM dataset. On average, our algorithm improves translation and rotation accuracy by 10.7% and 20.0%, respectively, compared to the best-performing baseline APDGICP. Figure 6 visualizes the estimated trajectories, where our approach (red lines) demonstrates superior accuracy in most splits, particularly in those involving large rotations.
As presented in Figure 5, we visualize the projection of consecutive point clouds from the loop3 split onto the corresponding images. With velocity compensation, the actual motion of dynamic points can be approximated, while the feature extraction algorithm selects keypoints that remain stationary and maintain strong continuity between frames.
Figure 7 illustrates the translation and rotation errors along the full cp split, excluding the two learning-based algorithms with large errors for clarity. While our method does not always yield the lowest translation error over the entire trajectory, it consistently achieves the best rotation accuracy. This outcome further validates the effectiveness of our invariant velocity preprocessing, SE(3)-equivariant network design, and adaptive loss function in robustly encoding and estimating rotational motion.
The comparative results on our self-collected dataset are also summarized in Table I and visualized in Figure 8. Due to the lower angular resolution of the employed 4D radar compared to that in the NTU4DRadLM dataset, all methods experience performance degradation. Nonetheless, our approach still attains the best translation and rotation accuracy across both splits, despite being trained solely on the NTU4DRadLM dataset, which demonstrates its strong generalization capability.
IV-D Ablation Study
To further validate the effectiveness of our proposed algorithm, we conduct an ablation study on the NTU4DRadLM dataset by creating three variant versions of Equi-RO:
  • Equi-RO with No Invariant Velocity Compensation (NIV): In this variant, node velocities are taken directly from the original Doppler values $v_{i}^{dop}$, and edge velocities are computed as $v_{i}^{dop}-v_{j}^{dop}$, rather than using the compensated counterparts $v_{i}^{dop^{\prime}}$ and $v_{ij}$.
  • Equi-RO with No Equivariant Network Design (NEN): Here, the LN-based layers are replaced by conventional Linear, ReLU, and Pooling layers, and the EGNN is substituted with a standard MPNN, removing the equivariance property from the network.
  • Equi-RO with No Adaptive Loss (NAL): In this case, the pitch loss $\mathcal{L}_{p}$ and yaw loss $\mathcal{L}_{y}$ are omitted, and fixed weights ($s_{r}=-8.0$, $s_{t}=-3.0$) are assigned to the rotation and translation losses instead of using learnable balancing parameters.
The quantitative results of this ablation study are summarized in Table II.
TABLE II: Ablation study results on the NTU4DRadLM dataset
@@ IV-D Ablation Study
NIV 5.16 0.0542 4.92 0.0601 8.12 0.0432 4.31 0.0088
NEN 9.60 0.1688 13.27 0.1960 8.97 0.0376 7.97 0.0101
NAL 5.32 0.0498 4.76 0.0744 9.11 0.0504 5.24 0.0222

### v2

@@ III-C Graph Construction
After pre-processing, the graph is constructed for each frame of point cloud by treating each point as a node. A cost function $f(i,j)$ is defined to select node pairs for edge generation, preferring neighboring points with similar velocities:
@@
where $\lambda$ is a weighting parameter. Edges are established between node $i$ and its $k$ nearest neighbors in the feature space.
Subsequently, to effectively fuse spatial and physical information, we organize the input into three distinct sets based on their transformation properties:
Geometric equivariant node feature set $\mathbf{X}\in\mathbb{R}^{N\times 3}$: This set contains Euclidean coordinates $\bm{\rho}_{i}$ of each node $i$ which reflect the geometric information and are inherently SE(3)-equivariant: $\mathbf{X}=\{\mathbf{X}_{i}\}=\{\bm{\rho}_{i}\}$.
Physical invariant node feature set $\mathbf{Y}\in\mathbb{R}^{N\times 3}$: This set represents the intrinsic properties of each node $i$ that remain stable under transformation, including the SO(3)-invariant compensated node velocity $v_{i}^{dop^{{}^{\prime}}}$, and SE(3)-invariant RCS value $r_{i}$ and the degree $d_{i}$, which is the number of edges connected to the node $i$: $\mathbf{Y}=\{\mathbf{Y}_{i}\}=\{[v_{i}^{dop^{{}^{\prime}}},r_{i},d_{i}]\}$.
Relative invariant edge feature set $\mathbf{Z}\in\mathbb{R}^{E\times 2}$ ($E$ is the total number of edges): To capture local relative information between nodes $i$ and $j$, the edge features consist of the compensated edge relative velocity $v_{ij}$, which remains SO(3)-invariant and further satisfies SE(3)-invariance for nodes with the same velocity, and the SE(3)-invariant Euclidean distance $\left\|\bm{\rho}_{i}-\bm{\rho}_{j}\right\|$: $\mathbf{Z}=\{\mathbf{Z}_{ij}\}=\{[v_{ij},\left\|\bm{\rho}_{i}-\bm{\rho}_{j}\right\|]\},j\in\mathcal{N}(i)$, where $\mathcal{N}(i)$ is the set of neighbors of $i$ in the graph.
These constructed feature sets ($\mathbf{X},\mathbf{Y},\mathbf{Z}$), and the graph structure $\mathbf{G}$, are built independently for both the source and target frames (denoted by subscripts $s$ and $t$ in Fig. 2) to serve as paired inputs for the subsequent network.
@@ III-D Network Design
Following graph construction and feature preparation, we establish a novel framework that processes the input equivariant and invariant features for both the source and target point clouds independently. The framework employs equivariant networks to handle different feature types and progressively integrate geometric structure with physical properties, yielding the respective final per-point feature set $\mathbf{F}_{s}$ and $\mathbf{F}_{t}$ for estimating the relative transformation between the two frames.
@@ III-D1 Feature Extraction
We employ a two-stage cascaded architecture for feature extraction of each frame. An SO(3)-equivariant Dynamic Graph Convolutional Neural Network (DGCNN) [25] first processes equivariant node feature set $\mathbf{X}$ of each frame of point clouds and produces SO(3)-equivariant output feature $\mathbf{X^{\prime}}$. The second module is an SE(3)-equivariant GNN (EGNN) module that, by integrating the processed node feature set $\mathbf{X^{\prime}}$ with the node and edge invariant feature set $\mathbf{Y}$ and $\mathbf{Z}$, outputs the final feature set $\mathbf{F}$.
The first DGCNN network aggregates features from each point and its neighbors by a local graph, generating richer local geometric representations. Before applying the network, we compensate the dominant translational component using ego-velocity and time interval, so that the SO(3)-equivariant layers can focus on rotation-consistent feature learning, and the whole system predicts the residual SE(3) motion on top of the translational initialization. A stack of LN [20] network layers is applied in the DGCNN framework. LN modifies traditional MLP layers such as Linear, ReLU and Pooling layers for equivariance on semi-simple Lie groups, and preserves SO(3)-equivariance in our case. After three iterations, the network outputs the processed node geometric feature set $\mathbf{X^{\prime}}$, whose dimension is lifted to $N\times(C^{\prime}\times 3)$ by LN and retains SO(3)-equivariance with SE(3)-aware pose reasoning.
The processed node feature set $\mathbf{X^{\prime}}$ of each frame is then passed to the second EGNN module together with physical invariant node feature set $\mathbf{Y}$ and relative invariant edge feature set $\mathbf{Z}$ of this frame for further graph feature aggregation. For each node $i$ and its neighbors $\mathcal{N}(i)$ connected by edges, the features are aggregated and updated as follows:
@@
$\mathbf{X}^{{}^{\prime}}_{i}$ $=\mathbf{X}^{{}^{\prime}}_{i}+\frac{1}{\left|\mathcal{N}(i)\right|}\underset{j\in\mathcal{N}(i)}{\sum}(\mathbf{X}^{{}^{\prime}}_{i}-\mathbf{X}^{{}^{\prime}}_{j})\phi_{E}(\mathbf{m}_{ij}),$ (11)
where $\mathbf{m}_{ij}$ represents the invariant message passed from node $j$ to $i$ under the message passing neural network framework [10], $\mathbf{X}^{{}^{\prime}}_{i}$ and $\mathbf{Y}_{i}$ are processed equivariant and invariant features of node $i$, respectively. $\phi_{I_{1}},\phi_{I_{2}},\phi_{E}$ are MLPs.
Differing from the original EGNN that uses $\left\|\mathbf{X}^{{}^{\prime}}_{i}-\mathbf{X}^{{}^{\prime}}_{j}\right\|^{2}$ for invariant feature extraction, inspired by [36], we employ a Gaussian kernel function $\exp(\gamma\left\|\mathbf{X}_{i}^{{}^{\prime}}-\mathbf{X}_{j}^{{}^{\prime}}\right\|)$ with a learnable parameter $\gamma$ to transform the SO(3)-equivariant features $\mathbf{X^{\prime}}$ into SE(3)-invariant representations.
After four iterations of the EGNN, the network outputs final node equivariant feature set $\mathbf{X}_{F}=\{\mathbf{X}^{{}^{\prime}}_{i}\}$ and node invariant feature set $\mathbf{Y}_{F}=\{\mathbf{Y}_{i}\}$ for i=1,2,,Ni=1,2,...,N for the processed frame. Each frame’s final per-point feature set $\mathbf{F}$ is defined as the concatenation of $\mathbf{X}_{F}$ and $\mathbf{Y}_{F}$.
@@ III-D2 Transformation Estimation
In this module, the $\mathbf{L}_{2}$ distances are computed between every feature vector $\mathbf{f}_{s,i}$ and $\mathbf{f}_{t,j}$ in the source and target feature sets $\mathbf{F}_{s}$ and $\mathbf{F}_{t}$, respectively, forming a pairwise-correspondence matrix. The $M$ pairs with the smallest feature distances are selected as keypoint correspondences. A learnable similarity metric is then defined as: $s_{ij}=\exp(-\beta(d_{ij}-\alpha))$, where $s_{ij}$ denotes the similarity between source point $i$ and target point $j$, $d_{ij}$ is their feature distance, and $\alpha,\beta$ are learnable parameters that enhance the robustness to outliers.
Given the similarity matrix $\mathbf{S}=\{s_{ij}\}_{M\times M}$, we obtain a soft assigning weight matrix $\mathbf{W}=\{w_{ij}\}_{M\times M}$ of correspondences via the differentiable Sinkhorn algorithm [29]. A weighted Singular Value Decomposition (SVD) finally estimates the transformation $(\bm{R},\bm{t})$ between the two frames.
@@ III-D3 Loss Function
The proposed framework is trained in an end-to-end manner using the ground truth transformation between frames as supervision. Specifically, we design a novel loss function that jointly considers rotation, translation, and orientation errors. The estimated rotation $\bm{R}$ is first converted to Euler angles $\mathbf{E}$, and the loss $\mathcal{L}$ between $(\mathbf{E},\mathbf{t})$ and the ground truth $(\mathbf{E}_{gt},\mathbf{t}_{gt})$ consists of four terms: rotation loss $\mathcal{L}_{r}$, translation loss $\mathcal{L}_{t}$, pitch loss $\mathcal{L}_{p}$, and yaw loss $\mathcal{L}_{y}$:
$\mathcal{L}_{r}$ $=\left\|\mathbf{E}-\mathbf{E}_{gt}\right\|,\mathcal{L}_{t}=\left\|\mathbf{t}-\mathbf{t}_{gt}\right\|,$
$\mathcal{L}_{p}$ $=\left\|\mathbf{E}^{pitch}-\mathbf{E}_{gt}^{pitch}\right\|,\mathcal{L}_{y}=\left\|\mathbf{E}^{yaw}-\mathbf{E}_{gt}^{yaw}\right\|,$ (12)
where $\mathbf{E}^{pitch}$ and $\mathbf{E}_{gt}^{pitch}$ denote the pitch angles, and $\mathbf{E}^{yaw}$ and $\mathbf{E}_{gt}^{yaw}$ denote the yaw angles, in the estimated and ground truth rotations, respectively. We include additional pitch loss $\mathcal{L}_{p}$ and yaw loss $\mathcal{L}_{y}$ for finer-grained rotation supervision. Geometrically, pitch and yaw directly affect forward motion, whereas roll primarily influences lateral and vertical components, which have less impact on forward trajectory error and can be sufficiently constrained by the main rotation loss $\mathcal{L}_{r}$.
@@ IV-A Datasets
Two datasets are used for evaluation: the NTU4DRadLM dataset [34] and a self-collected dataset acquired using a vehicle-mounted 4D radar and high-precision GPS.
@@ IV-A1 NTU4DRadLM Dataset
This dataset is designed for 4D radar-based localization and mapping. It provides ground-truth odometry and includes sequences from both low-speed robot platforms and high-speed vehicles. It was collected using the Oculii Eagle 4D radar, which has $0.5^{\circ}$ azimuth and $1^{\circ}$ elevation resolution at 12 Hz operating frequency. We use its enhanced radar point cloud and select loop1 for training, loop2 for validation, and the remaining four splits for testing.
@@ IV-A2 Self-Collected Dataset
To further assess generalization, we constructed a campus dataset using the Geometrical Pal R7861B 4D radar. As shown in Fig. 4, the radar was mounted on an SUV equipped with a high-precision GPS system to provide ground-truth odometry. The radar has $1.9^{\circ}$ azimuth and $3.5^{\circ}$ elevation resolution, operates at 15 Hz. All learning-based models trained on NTU4DRadLM are directly tested on this dataset to compare cross-dataset generalization.
@@
Figure 5: Projection of adjacent point clouds onto the corresponding images from loop3 split, where point colors represent compensated node velocities and white circles denote selected keypoints used for matching.
Figure 6: Comparative odometry results of Equi-RO and baselines on the NTU4DRadLM dataset for the cp (a), garden (b), nyl (c), and loop3 (d) splits.
@@
Figure 7: Comparative odometry results of Equi-RO and baselines on the two splits of self-collected dataset.
Figure 8: Comparative odometry results on large-rotation segments of cp (a), garden (b), nyl (c), and loop3 (d) splits of the NTU4DRadLM dataset.
TABLE II: Quantitative results on large-rotation segments of the NTU4DRadLM dataset
@@
ICP 70.04 0.6056 20.59 0.2884 63.04 0.2001 53.81 0.5660
NDT Failed Failed 5.76 0.1337 7.23 0.0881 6.84 0.5276
GICP 16.51 0.5375 4.73 0.1706 8.44 0.0816 6.68 0.5043
APDGICP 17.61 0.5189 3.63 0.1578 6.83 0.0846 4.94 0.4480
KISS-ICP 14.35 0.4219 5.33 0.1785 7.23 0.0982 7.23 0.5150
Radar4Motion 13.26 0.4398 4.86 0.1813 7.38 0.0738 6.52 0.5109
RaFlow 50.43 2.1401 43.78 0.5364 37.14 0.2808 60.96 0.9989
CAO-RONet 29.30 1.9730 50.45 0.2681 28.83 0.3657 72.31 1.6608
Equi-RO 7.23 0.2588 3.38 0.1457 6.47 0.0762 4.64 0.4405
Figure 10: Rotation (left) and translation (right) errors under different dynamic-point ratios for Equi-RO and the ablation variant I. Points with absolute compensated Doppler velocities greater than 0.1 m/s are treated as dynamic, and numbers on top of bars denote frame counts.
IV-B Experimental Setup
IV-B1 Training Details
We train Equi-RO on a single NVIDIA A100 GPU using Adam with a weight decay of $10^{-3}$ and a batch size of 4 for 50 epochs. The initial learning rate is $10^{-4}$ and decays by a factor of 0.1 every 10 epochs. For hyperparameters, we set the weight $\lambda=1$ and the number of nearest neighbors $k=6$ in graph construction.
IV-B2 Compared Algorithms
We first select several traditional registration methods ICP, NDT, and GICP as baselines. We also evaluate an ICP-based method KISS-ICP [31], a 4D radar odometry method Radar4Motion [16], and Adaptive Probability Distribution-GICP (APDGICP) [35], the State-of-the-Art (SOTA) 4D radar odometry method on NTU4DRadLM dataset. For learning-based methods, we include RaFlow [7] and CAO-RONet [19]. These two methods are also trained on the loop1 split of NTU4DRadLM using their default settings.
IV-B3 Metrics
We adopt the evo library [11] to compute the relative translation error ($t_{rel}$) and relative rotation error ($r_{rel}$) for each algorithm, by comparing registered trajectories with the ground truth. We average errors over 100 equal segments (with $\delta=L/100$) of each trajectory.
IV-C Experimental Results
Fig. 5 qualitatively shows that velocity compensation can help estimate the actual motion of dynamic points, while the feature extraction can select keypoints that remain stationary and consistent between frames.
We present quantitative comparison of our Equi-RO against baseline algorithms in Table I. In the NTU4DRadLM dataset, our method achieves the lowest translation and rotation errors on the cp, garden, and loop3 splits, and ranks second best rotation error on the nyl split. GICP, APDGICP, KISS-ICP and Radar4Motion attain competitive results, while learning-based approaches RaFlow and CAO-RONet exhibit inferior performance. This discrepancy is likely because they are originally developed for the earlier VoD dataset [24], and appear less suitable to the newer NTU4DRadLM dataset. On average, our algorithm improves translation and rotation accuracy by 10.7% and 13.4%, respectively, compared to the best baselines. Fig. 6 visualizes the estimated trajectories, where our method (red lines) demonstrates superior accuracy in most splits.
The comparative results on self-collected dataset are also shown in Table I and Fig. 7. All methods degrade on this dataset with lower radar resolution. Nonetheless, our method still attains the best translation and rotation accuracy across both splits, despite being trained solely on the NTU4DRadLM dataset, demonstrating its strong generalization capability.
We further evaluate performance on sharp-turn segments from the NTU4DRadLM dataset. Equi-RO consistently achieves the best or second-best accuracy, as shown in Fig. 8 and Table II. These results confirm that the proposed equivariant and invariant features, and the equivariant network design are especially effective in scenarios with large rotations.
IV-D Ablation Study
We conduct an ablation study on the NTU4DRadLM dataset by creating five variants of Equi-RO:
(I) w/o invariant velocity compensation: In this variant, node velocities are set to the original Doppler values $v_{i}^{dop}$, and edge velocities are computed as $v_{i}^{dop}-v_{j}^{dop}$, rather than using the compensated counterparts $v_{i}^{dop^{\prime}}$ and $v_{ij}$.
(II) w/o equivariant network design: Here, the LN-based layers are replaced by conventional Linear, ReLU, and Pooling layers, and the EGNN is substituted with a standard message passing neural network, removing the equivariance property.
(III) w/o adaptive loss: In this case, the pitch loss $\mathcal{L}_{p}$ and yaw loss $\mathcal{L}_{y}$ are omitted, and fixed weights ($s_{r}=-8.0$, $s_{t}=-3.0$) are assigned to the rotation and translation losses instead of learnable parameters.
(IV) w/o Doppler velocity: The compensated node velocity $v_{i}^{dop^{\prime}}$ and edge velocity $v_{ij}$ are set to zero in this case.
(V) w/o initial translation: In this ablation, we remove the initial translation based on ego-velocity.
The quantitative results are summarized in Table III.
TABLE III: Ablation study results on the NTU4DRadLM dataset
Method cp garden nyl loop3
I 5.16 0.0542 4.92 0.0601 8.12 0.0432 4.31 0.0088
II 9.60 0.1688 13.27 0.1960 8.97 0.0376 7.97 0.0101
III 5.32 0.0498 4.76 0.0744 9.11 0.0504 5.24 0.0222
IV 5.25 0.0523 4.94 0.0614 8.67 0.0505 6.78 0.0109
V 5.21 0.0514 3.97 0.0427 6.87 0.0401 4.05 0.0095